Hi! Once again, I'm asking help for one of these pesky (for me, at least) money problems. The answer is supposed to be $2,000, but the answer that I'm always getting ends up to be way off the mark. The question is as follows:
A man had $6,000. He spent part of it to pay for a jeep and invested the remaining amount as follows: 1/3 of it at 6% and 2/3 of it at 4% simple interest for 1 year. The income from the two investments was $140. What was the cost of the jeep?
Thanks for the time and help! :)

Found 2 solutions by scott8148, richwmiller:
Answer by scott8148(6628)

[(6000 - j)(1 / 3)(.06)] + [(6000 - j)(2 / 3)(.04)] = 140

(6000 - j)(.14 / 3) = 140

6000 - j = 3000 ___ 3000 = j


not 2000 ___ sorry...

Answer by richwmiller(17219)

j= cost of jeep
x=amt at 6%
y=amt at 4%
x+y+j=6000
x+y=6000-j
x=1/3(6000-j)
y=2/3(6000-j)
.06*x+.04y=140
j = 3000, x = 1000, y = 2000